trentin:

1 over 2 x + 3 = −x + 7

trentin:

\[-\frac{ 1 }{ 2 }x+3=-x+7\]

ThisGirlPretty:

@Bearclaws72

trentin:

it x=8

Ultrilliam:

I got \(x=\frac{8}{3}\)

trentin:

how did you get \[\frac{ 8 }{ 3 }\]

trentin:

@Ultrilliam

Ultrilliam:

Don't trust me though, I'm not great at math :|